### v1

# Improving Concept Alignment in Vision-Language Concept Bottleneck Models

Abstract
Concept Bottleneck Models (CBM) map the input image to a high-level human-understandable concept space and then make class predictions based on these concepts. Recent approaches automate the construction of CBM by prompting Large Language Models (LLM) to generate text concepts and then use Vision Language Models (VLM) to obtain concept scores to train a CBM. However, it is desired to build CBMs with concepts defined by human experts instead of LLM generated concepts to make them more trustworthy. In this work, we take a closer inspection on the faithfulness of VLM concept scores for such expert-defined concepts in domains like fine-grain bird species classification and animal classification. Our investigations reveal that frozen VLMs, like CLIP, struggle to correctly associate a concept to the corresponding visual input despite achieving a high classification performance. To address this, we propose a novel Contrastive Semi-Supervised (CSS) learning method which uses a few labeled concept examples to improve concept alignment (activate truthful visual concepts) in CLIP model. Extensive experiments on three benchmark datasets show that our approach substantially increases the concept accuracy and classification accuracy, yet requires only a fraction of the human-annotated concept labels. To further improve the classification performance, we also introduce a new class-level intervention procedure for fine-grain classification problems that identifies the confounding classes and intervenes their concept space to reduce errors. Source codes are available at https://github.com/NMS05/Improving-Concept-Alignment-in-Vision-Language-Concept-Bottleneck-Models.
Keywords:
Concept Bottleneck Model Interpretability Semi-Supervised
1 Introduction
Concept Bottleneck Models (CBM) [14] are a popular choice for building interpretable models with high-level concepts. These models first map the input image to a low-dimensional (hence the term “bottleneck”) expert-defined concept space where each concept is readily understandable by a human. For example, the bird species ‘Vermilion Flycatcher’ can have concepts like ‘red-body’, ‘black-wing’, and ‘red-crown’, etc. A linear classifier then makes class predictions on top of these concepts. There are two fundamental requirements to build a CBM - a well-defined non-ambiguous concept set and the corresponding concept labels for every sample. Recent approaches [25, 46] automate the above steps and scale CBM to larger datasets (like ImageNet) by leveraging Large Language Models (LLM) and Vision Language Models (VLM). They typically involve generating an initial concept set by prompting a LLM with few shot examples and then extracting a sub-set of those concepts that are conducive for discrimination (classification). When an image and this filtered concept set  [^1] The number of LLM generated concepts are usually an order of magnitude larger than the expert-defined concepts. is passed as a input to the contrastively pre-trained VLM, it’s image-text alignment scores act as “concept scores or concept labels” for that sample. These concept labels can be used to train another CBM classifier [25] or can be directly used to predict the classes [46], thus turning a VLM into CBM (referred to as VL-CBM from here on).
While LLMs possess rich world knowledge [27, 11, 34] and can generate a potential (large) concept set across any domain, there are certain instances where an expert-defined concept is preferred.
  • A small and concise set of expert-defined bottleneck concepts enables easier error analysis and intervention, which is crucial for interpretable models.
  • In medical domains, such as Hematology [36, 2], it is desired to use the expert-defined concept set rather than that generated from LLMs, as the former comes from the more trusted pathologists.
  • For faithful “visual” classification, all the concepts must be visually groundable, which is guaranteed with expert-defined concept set [38]. In comparison, for problems like bird species classification, LLMs can generate concepts like “migratory birds” that are not salient for visual classification.
In such cases with limited yet preferable expert concepts, do the VLM concept scores faithfully represent the visual input and are they efficacious to build VL-CBMs with desired performance? To answer this question, we take a closer inspection on the CLIP [28] concept scores for expert-defined concepts on classification datasets like CUB [38], RIVAL [21], and AwA2 [42]. Ideally, it is desired for the CLIP model to precisely associate a visual concept (appearance, color, shape, etc.) to the corresponding visual input. Our investigations expose two problems,
  • Low Concept Accuracy. The CLIP (VLM) model has a low concept accuracy despite the VL-CBM achieving a high classification performance, which makes the resulting model less faithful.
  • Incorrect Concept Association. For challenging classification problems like CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts. It struggles to correctly attribute the fine-grain concepts to the visual input.
Hence, to improve the reliability of VL-CBM, there is a need to improve concept alignment i.e., activate the desired/truthful concepts for a given image, which requires supervision. But obtaining the supervisory concept labels for all training samples is cumbersome. To solve this problem, we propose a novel Contrastive Semi-Supervised (CSS) learning approach that can improve the concept alignment in VL-CBM with fewer labeled concept examples. Our approach encourages consistent concept scores for intra-class samples whilst discriminating (contrasting) them from the other classes. It then uses a few labelled concept examples (semi-supervision) per class to align them with the ground truth. Extensive experiments on three benchmark datasets CUB [38], RIVAL [21], and AwA2 [42] show that our CSS method substantially increases the concept accuracy (+39.1% for CUB, +18.63% for RIVAL, +31.11% for AwA2) and enhances the overall classification accuracy (+5.61% for CUB, +2.84% for RIVAL, +3.06% for AwA2) with only a handful of human-annotated concept labels per class. We also validate our CSS approach in a medical domain using the blood cell attribute dataset WBCAtt [36].
Intervention is the fundamental advantage of a CBM over other interpretable models - if a classifier makes an incorrect prediction, one can inspect the concept scores and rectify the problem, so that the model makes the desired prediction. Intervention in CBM is typically performed at an instance level [14, 25], where a handful of error images are arbitrarily chosen for debugging. But, choosing the error images are non-trivial for fine-grain multi-class classification problems which have fewer samples per class and even visually similar classes. In this work, we propose a class-level intervention procedure for fine-grain classification problems, where we first compute an error matrix to identify the “confounding classes” (visually similar but semantically different) and then intervene the error images of these classes such that it reduces the total error for the confounding classes whilst increasing the overall classification accuracy. The primary contributions of this paper are summarized below,
  • Validating the faithfulness of VL-CBM. We investigate the concept scores for expert concepts and show that CLIP models have poor concept alignment and they struggle to correctly associate the fine-grain concepts to visual input.
  • Improving concept alignment. To address the above issue and to improve the reliability of VL-CBM, we propose a novel Contrastive Semi-Supervised (CSS) approach which uses a fraction of the human-annotated concept labels to increase the concept accuracy and classification accuracy.
  • Class-level Intervention. We propose an intervention procedure for fine-grain classification problems which reduces the total error for confounding classes and further increases the classification performance.
2 Related Work
Post-hoc interpretable models. Post-hoc methods [30, 29, 17, 4, 8] convert a pre-trained black-box classifier into an interpretable model using saliency maps. They highlight the regions of the input space [31] that influence a classifiers’ prediction and it helps one understand the salient features of a particular class. Though saliency maps can reveal the semantic features (or spurious correlations) that a classifier focuses on, their interpretation can be subjective and may require domain expertise.
Prototype-based and Concept-based interpretable models. In stark contrast to post-hoc methods, these models can be trained and are interpretable by design. Prototype-based methods [6, 24, 32, 37, 7] learn interpretable visual feature prototypes for every class from the training samples. During inference, the distance (similarity score) between the test sample feature and the learned prototypes determine the classifiers decision. BotCL [39] is a prototype-based method that is closely related to our work. It applies contrastive learning to the prototype space of a vision-only model, while we focus on contrastive + semi-supervised learning on the concept space of a vision-language (multimodal) model. Concept-based methods [12, 14, 48] map the input to a high-level concept space defined by human experts, where these concepts are readily interpretable. The classifiers decision is based on the predicted concept scores. The main advantage of concept-based methods is that they allow test-time intervention [14, 12, 19, 1, 5] and they are easier to debug compared to other methods.
VLM based interpretable models (VL-CBM). The problem with concept-based models is that they require concept labels for every training sample [14] or a lot of training images representing each of those concepts [12, 48]. To solve this problem, recent approaches [25, 46, 49, 45, 41, 22, 23] leverage contrastively pre-trained Vision-Language Models, like CLIP [28], to automatically generate concept labels from their image-text alignment scores. These models have interpretable natural language concepts [18, 44] that are typically generated by a LLM and this automated approach scales them to larger datasets.
Enhancing faithfulness of VL-CBMs. Several recent works have focused on addressing the shortcomings of VLM based concept bottlenecks to improve their reliability. They constrain the class predictions to focus on a sub-set of salient concepts [40, 33, 26], make concept scores robust to perturbations [15], and obtain reliable concept score distributions by modeling uncertainties [13] and by composing energy functions [43]. In our work, we enhance the faithfulness of VL-CBM by using human (expert) preferred concepts and improve the concept scores of VLM in a semi-supervised setup using a few concept labels.
3 Methodology
We formally define VL-CBM in Sec. 3.1 and investigate the faithfulness of CLIP concept scores in Sec. 3.2. We then propose our Contrastive Semi-Supervised (CSS) method to improve concept alignment in Sec. 3.3.
3.1 Vision-Language Concept Bottleneck Models
Consider a dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the images, $y_{i}$ denotes the labels for $k$ classes, and $N$ is the total number of training samples. While a black-box classifier $f:x\rightarrow\mathbb{R}^{k}$ directly maps an image $x$ to its label $y$, a CBM consists of a backbone $g:x\rightarrow\mathbb{R}^{c}$ which first maps the image $x$ to a concept space consisting of $c$ pre-defined concepts and a linear classifier $h:\mathbb{R}^{c}\rightarrow\mathbb{R}^{k}$ that maps the predicted concepts to its corresponding label $y$. These concepts represent the salient visual features that are crucial for solving the classification task.
However, training the backbone $g$ requires $N\times c$ concept labels which is tedious to obtain. VL-CBM overcome this issue by leveraging the image-text alignment scores of contrastively pre-trained Vision-Language Models (like CLIP) as concept labels. Concretely, let $T=\{t_{1},t_{2},...,t_{c}\}$ be the set of LLM generated or expert-defined text concepts. Now the $c$ concept labels for an image $x_{i}$, which is denoted as $C_{i}$, is obtained by $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}$, where $\cdot$ denotes the dot product, $E_{I}$ and $E_{T}$ denotes the image and text encoder of the CLIP model that maps the images and text concepts into a shared feature space. Now, with the availability of image $x_{i}$, concept label $C_{i}$, and class label $y_{i}$ for $N$ samples, the backbone $g$ and the classifier $h$ can be independently or jointly trained.
Table 1: Evaluation of CLIP concept scores and VL-CBM classification performance.
   CUB    RIVAL    AwA2 Classification Accuracy (%)    75.84    95.63    90.14 Concept Accuracy (%)    24.43    58.85    49.02 No.of Concepts    312    18    85 Difficulty of Concepts    Hard    Easy    Medium
3.2 Faithfulness of Vision-Language Concept Bottleneck Models
We validate the faithfulness of VL-CBM for expert concepts across three image classification datasets namely Caltech-UCSD Birds (CUB) [38], Rich Visual Attributes with Localization (RIVAL) [21], and Animal with Attributes (AwA2) [42]. These datasets provide expert concepts and the corresponding ground-truth concept labels required for our evaluation. We input the images and the expert (text) concepts to a frozen CLIP model to obtain the concept scores as described in Sec. 3.1. Our objective here is to evaluate the quality and faithfulness of the VLM concept scores by comparing them against the ground-truth concept labels and evaluate their classification performance with a linear classifier. The concept evaluation results are presented in Table 1. Our investigations expose two problems with CLIP concept scores.
Low Concept Accuracy. From the table, the CLIP model has a low concept accuracy despite the VL-CBM achieving a high classification accuracy. This implies that a higher classification performance does not guarantee a faithful concept understanding, which makes the VL-CBM less reliable. Moreover, it becomes difficult to intervene such a model during test-time, in order to rectify prediction errors.
Incorrect Concept Association. A closer inspection of these concepts in Fig. 2 (CLIP Concept Scores) reveal that, for fine-grained concepts like in CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts. The untrained model struggles to correctly associate the color/texture concepts to the corresponding bird part, which is crucial for bird species classification.
The lower concept accuracy and the errors in concept association is a result of the poor relational and compositional understanding of CLIP models i.e., they make errors in associating objects to their attributes [47], especially in a multi-object and multi-attribute setting [16]. Hence, to improve the reliability and faithfulness of VL-CBM, there is a need to improve concept alignment i.e. activate the truthful concepts for the given image.
Figure 1: (a) Architecture of VL-CBM with concept projection (bootstrap). (b) Overview of the Contrastive Semi-Supervised (CSS) learning method.
3.3 Improving concept alignment in VL-CBM
We propose a novel Contrastive Semi-Supervised (CSS) learning method that can improve the concept alignment in VL-CBM with a few labelled concept examples per class. The overall framework is illustrated in Fig. 1.
From Table 1, we can see that the frozen CLIP model achieves a low yet non-trivial concept accuracy. We bootstrap this by adding a learnable linear concept projection layer (see Fig. 1a) that enhances the CLIP concept scores by directly predicting the $c$ concepts from the average pooled visual patch tokens as,
$C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}+LN(E_{I}^{*}(x_{i}))\cdot W_{CP},$ (1)
where $E_{I}(x_{i})\in\mathbb{R}^{1\times 512}$ is the image feature, $E_{T}(T)\in\mathbb{R}^{c\times 512}$ is the text feature for $c$ concepts, $E_{I}^{*}(x_{i})\in\mathbb{R}^{1\times 768}$ denotes the average pooled patch tokens from the visual encoder, $W_{CP}\in\mathbb{R}^{768\times c}$ is the learnable concept projection layer and $LN$ denotes layer normalization. From the concept scores $C_{i}\in\mathbb{R}^{1\times c}$, a linear classifier $W_{K}\in\mathbb{R}^{c\times k}$ makes class predictions $K_{i}=C_{i}\cdot W_{K}$ for the image $x_{i}$, where $K_{i}\in\mathbb{R}^{1\times k}$ denotes the class logits for the $k$ classes.
Now the model must be optimized to improve concept alignment and achieve high classification accuracy. To improve concept alignment, the predicted concepts can be compared with the ground truth concept labels. However, the availability of fewer concept labels (semi-supervision) makes it less effective in improving concept alignment across other unlabelled samples. To mitigate this, we augment a contrastive learning objective to the concept space with the following intuition. The contrastive objective encourages “consistent” concept scores for intra-class samples and the concept labels for a few of these samples (semi-supervision) guide or align the predicted concepts with the ground truth. Furthermore, contrastive learning discriminates the concept scores of inter-class samples thus aiding in classification. Such a learning objective of enforcing consistency between sample pairs guided by a few labeled examples [3, 20] has not been studied in VL-CBMs before.
To achieve this (see Fig. 1b), we perform pairwise sampling such that each training sample consists of a pair of images $(x_{i},x_{j})$ from the same class and every other sample in the mini-batch belongs to different classes. There are $n$ samples ($2n$ images) in each mini-batch. For these image pairs, we compute the concept scores $(C_{i},C_{j})$ and class predictions $K=[K_{i},K_{j}]~{}$ [^2] The square brackets denote the tensors are stacked across batch dimension and no longer need to be in pairs.. Let $g=[g_{i},g_{j}]$ and $y=[y_{i},y_{j}]$ be the corresponding ground-truth concept labels and class labels. We jointly train the concept projection layer and the linear classifier to minimize the following loss objective.
[…]
$L=L_{contrastive}+L_{CE}+L_{concept},$ (2)
$L_{contrastive}=\frac{1}{n}\sum_{i,j\in n}-\log\frac{e^{sim(C_{i},C_{j})/\tau}}{\sum_{m\neq i}e^{sim(C_{i},C_{m})/\tau}},$
$L_{CE}=\frac{1}{2n}\sum_{l\in 2n}L_{ce}(K_{l},y_{l}),$
$L_{concept}=\frac{1}{2n}\sum_{l\in 2n}L1(\gamma\cdot(s(C_{l})-s(G_{l}))),\ \ \textrm{if}\ \ |G_{l}|\neq 0,$
$L_{concept}=0,\ \ \textrm{if}\ \ |G_{l}|=0,$
where $sim$ is the cosine similarity metric, $\tau$ is the temperature parameter, $L_{ce}$ and $L1$ are the standard cross entropy loss and mean absolute error, $s$ is the softmax normalization and $\gamma$ is a constant that increases the magnitude of the $L1$ loss function. $l$ denotes the index of the concept score, class logit, and label tensor within the mini-batch, after stacking along the batch dimension.
  • The contrastive loss ($L_{contrastive}$) optimizes $W_{CP}$ and encourages the concept scores of the same classes (positive pairs) to be consistent and discriminates it from the other classes (negative pairs).
  • The cross entropy loss ($L_{CE}$) trains the linear classifier $W_{K}$ to make class predictions from the concept scores.
  • The concept loss ($L_{concept}$) optimizes $W_{CP}$ to align the concept predictions with the ground truth concept labels and it primarily improves the concept accuracy. The concept loss is enabled if the concept labels are available ($|G_{l}|\neq 0$) or else it becomes zero.
4 Experiments
4.1 Experimental Setup
Dataset. We validate the effectiveness of our Contrastive Semi-Supervised (CSS) method in improving concept alignment on the CUB [38], RIVAL [21], AwA2 [42] datasets, and in a medical setting on the White Blood Cell Attribute (WBCAtt) dataset [36]. All the above datasets provide an expert-defined concept set and the corresponding ground-truth concept labels, which is crucial for our evaluation.
Model and Training. We use the PLIP ViT-B-16 [9] model available in Hugging Face for experiments on WBCAtt dataset and use the OpenCLIP ViT-B-16 model [10] for experiments on all other datasets. We use Adam optimizer to train all our models and evaluate them with classification accuracy and/or concept accuracy metrics.
Table 2: Improving Concept Alignment in VL-CBM using Contrastive Semi-Supervised (CSS) learning. We report the classification, concept, and WBC attribute accuracies (in %) as Class, Concept, and Attribute. The best performance is bolded, and the second-best is underlined. Note that WBCAtt uses the PLIP model.
Model CUB RIVAL AwA2 WBCAtt  Class  Concept  Class  Concept  Class  Concept Attribute ViT-B/16 (Black Box)  79.21 NA  99.47 NA   94 NA  66.11 CLIP VL-CBM [25, 46]  75.84 24.43  95.63 58.85  90.14 49.02 42.56 CSS VL-CBM (Ours)  81.45 63.53  98.47 77.48   93.2  81.13  64.19
4.2 Improving Concept Alignment
We empirically demonstrate the effectiveness of our Contrastive Semi-Supervised (CSS) approach in improving concept alignment for the CUB, RIVAL, AwA2, and WBCAtt datasets in Table 2.
Bird and Animal classification. We validate our CSS method with concept accuracy and classification accuracy metrics for CUB, RIVAL, and AwA2 datasets. We first train a black-box image classifier with frozen CLIP vision encoder (ViT-B/16) as the backbone to establish the baseline classification performance. The black-box classifier directly predicts the classes from the image features without any concept bottlenecks. We then train a CLIP VL-CBM classifier with expert-defined text concepts as described in Sec. 3.1. From the table, we can see that the classification performance of CLIP VL-CBM is on-par with the black-box classifier, however, it has a lower concept accuracy.
Figure 2: Visualization of $top-8$ concept scores for the CUB (birds) and AwA2 (animals) datasets. Incorrectly activated concepts are highlighted in red.
Hence, to improve the faithfulness and reliability of VL-CBM, we apply the concept projection later (bootstrap) and optimize the model with our CSS objective (CSS VL-CBM). Our CSS method not only substantially increases the concept accuracy, it also closes the gap with black-box methods on image classification performance. Especially for the CUB dataset, our CSS method surpasses the black-box classifier. The “number of concept labels per class” used to train the CSS VL-CBM model for the above datasets are as CUB - 9 labels (30%), RIVAL - 8 labels (0.1%), and AwA2 - 10 labels (1.4%). The values within brackets denote the “concept labels used for training as a percentage of the total concept labels provided by the dataset”. The results show that even with such small percentages of concept labels, our CSS method can substantially increase the concept accuracy. The enhanced and faithful concept scores after improving concept alignment with CSS is presented in Fig. 2 (CSS Concept Scores).
WBC Attribute prediction. WBCAtt contains White Blood Cell (WBC) images annotated with 31 “cell morphology” attributes that aids in diagnosis in Hematology. We directly evaluate the models on their attribute prediction accuracy and do not perform any image classification. While contemporary approaches [48, 46] use the HAM10000 [35] dataset to study VL-CBM in the medical domain, the availability of attribute labels in WBCAtt makes the latter suitable for our work. For our evaluation, we train the PLIP model [9] (a VLM model for Pathology AI), with 50% of the attribute labels. From the table, we can see that our CSS method substantially increases the attribute prediction accuracy and closes the gap with black-box attribute predictor.
Figure 3: Number of concept labels vs CSS VL-CBM performance.
Impact of the number of concept labels. Since our approach is semi-supervised, we further investigate the impact of the number of human-annotated concept labels on the CSS VL-CBM performance in Fig. 3.
For CUB and RIVAL datasets we progressively increase the concept labels used for training, for which we observe a proportional increase in the classification and concept accuracies.
For WBCAtt, we investigate how the black-box (PLIP vision encoder) attribute predictor compares against the CSS (PLIP) VL-CBM attribute predictor, for an increase in the attribute labels used for training. Note that there is no image classification task for WBCAtt. We can see that our CSS method outperforms black-box attribute predictor in a low data-resource setting.
4.3 Comparison with other methods
CUB. The CUB dataset is extensively used by the Explainable AI community to build and study concept-based interpretable models. Hence, we comprehensively compare our Contrastive Semi-Supervised VL-CBM method with contemporary VLM-based interpretable models on the CUB dataset. The results are presented in Table 3. For a fair comparison with LF-CBM [25] and PCBM [48], we perform two-stage training. In stage 1, a ResNet18 backbone with a linear concept predictor is trained with the concept labels generated by our CSS VL-CBM model. It minimizes the $L_{concept}$ loss function. In stage 2, a sparse-linear SGD classifier from scikit-learn is trained with 5000 steps as described in [48]. The backbone, concept predictor, and sparse-linear classifier together constitute a CBM.
Table 3: Comparison of CSS VL-CBM with contemporary methods on CUB.
Method Base CBM Training Classifier Class (%) Concept (%) LF-CBM [25] ResNet-18 2-stage Sparse-Linear 74.31 - PCBM [48] ResNet-18 2-stage Sparse-Linear 59.6 - CLIP VL-CBM ResNet-18 2-stage Sparse-Linear 68.77 16.53 CSS VL-CBM (Ours) ResNet-18 2-stage Sparse-Linear 72.85 62.8 CLIP VL-CBM ResNet-18 joint Linear 74.91 18.84 CSS VL-CBM (Ours) ResNet-18 joint Linear 80.36 63.17 CompDL [49] CLIP ViT-B/32 - Linear 52.6 - LaBo (full shot) [46] CLIP ViT-B/32 - Linear 81.0 - CSS VL-CBM (Ours) CLIP ViT-B/16 joint Linear 81.45 63.53
From the table, we can see that the classification accuracy of ResNet-18 CBM supervised by our CSS VL-CBM approach is on par with LF-CBM and achieves much higher performance than PCBM. However, we observe the best performance with joint training of the backbone, concept predictor, and linear classifier layers with $L_{concept}$ and $L_{ce}$ losses. Grad-CAM [31] visualization of the concepts learned by the (jointly trained) ResNet18 CBM supervised by our CSS VL-CBM is presented in Fig. 4. From the figure, we can see that the improved concept knowledge (activating the truthful concepts for a given image) of our CSS VL-CBM model was effectively transferred to the ResNet model. It is able to correctly associate a color or texture concept to the corresponding bird part. Also, our method performs better than LaBo [46] with just $312$ (CUB) expert concepts, while LaBo utilizes $10,000$ LLM-generated bottleneck concepts.
Figure 4: Grad-CAM visualization of the learned concepts.
Table 4: Comparison of CSS VL-CBM with other methods on RIVAL and AwA2.
Dataset RIVAL AwA2 CBM Method  Text2Concept   CSS  PCBM   ECBM   CSS Classification Accuracy (%) 95.38   98.86  88   91.2   93.2 Concept Accuracy (%) -   72.84  71.9   85.84   81.3
RIVAL and AwA2. We further compare CSS VL-CBM with contemporary methods on the animal classification datasets in Table 4. For RIVAL dataset, CSS achieves a higher classification performance compared to Text2Concept [22]. For AwA2 dataset, CSS achieves a higher classification accuracy compared to other methods, but achieves a slightly lower concept accuracy compared to ECBM. It is important to note that PCBM [13] and ECBM [43] do not use “text” concepts and they optimize the feature extractor (vision backbone) during training.
5 Intervention
When a CBM makes an error, one can inspect the incorrect predictions and manually tune the classifier weights or modify the concepts to fix those errors. This approach is called “intervention” or “model editing”. Conventional approaches to intervention are typically done at the instance-level [25, 48, 14], where a handful of error images are arbitrarily chosen for intervention. However, for fine-grain multi-class classification problems like CUB which have more classes (some bird species are visually similar) and relatively fewer samples per class, choosing the error images is non-trivial. Moreover, it is desired to perform intervention such that it increases the performance globally [48] instead of rectifying the errors for a few images. Hence, we propose a class-level intervention procedure where we first identify the confounding classes and then choose the error images of these classes for intervention. Confounding classes are bird species, that are visually similar yet semantically different and the model makes the most errors for these challenging classes. This class-level intervention is done in the following steps.
Figure 5: Confounding classes - birds that are visually similar yet belong to different sub-species.
Table 5: Class-level intervention of CSS VL-CBM. The confounding classes are as California Gull (CG), Western Gull (WG), Common Tern (CT), and Artic Tern (AT). The value within brackets denote the total number of test samples for that class.
Before Interv.   After Interv. Total error for CG (30) 27 14 Total error for WG (30) 15 9 CG misclassified as WG 13 4 WG misclassified as CG 3 1 Total error for CT (30) 21 11 Total error for AT (29) 3 2 CT misclassified as AT 12 6 AT misclassified as CT 0 0 Overall Classification Accuracy (%) 81.45 82.57
Step 1: Identify the confounding classes. First, we compute a $K\times K$ error matrix by evaluating the trained CSS VL-CBM model on the test set, where $K$ is the number of classes. The error matrix gives the error distribution for every class and the index of the matrix element with the highest value (error) gives the confounding class pairs. In our experiments, we identified two pairs of confounding classes “California Gull - Western Gull” and “Common Tern - Artic Tern” as shown in Fig. 5. The total number of classification errors for these confounding classes before intervention are presented in Table 5.
Step 2: Inspecting the concept space. Next, we inspect the concept scores of the error images in these confounding class pairs. We find that the concept score distribution of the confounding classes are nearly identical, which makes it difficult to discriminate them. Our model suffers from Type-2 error [25] i.e., the number of bottleneck concepts and their score distributions are insufficient to effectively discriminate the samples.
Step 3: Expanding the concept set. To mitigate the Type-2 error and to improve the overall performance, we expand the concept set by leveraging the concepts provided by LaBo [46]. Note that the new concepts to be added should be visually groundable by the CLIP model and hence we use the “filtered” concept set of LaBo. We choose $top-32$ concepts per confounding class and expand the concept set from the original 312 concepts to 440 concepts. Formally, let $T^{\prime}$ be the new set of 128 text concepts and $C^{\prime}_{i}=E_{I}(x_{i})\cdot E_{T}(T^{\prime})^{\intercal}$ be the CLIP concept scores for it. Now the concept scores for all the 440 concepts for an image $x_{i}$ is given by $C^{\prime}=$ Concat($C_{i},C^{\prime}_{i}$). During this intervention procedure we do not train a new concept projection layer for the expanded concept set.
Step 4: Training the classifier. We add a new linear classifier $W^{\prime}_{K}\in\mathbb{R}^{128\times 4}$ that directly predicts the confounding classes as $K^{\prime}=C^{\prime}\cdot W^{\prime}_{K}$. Then we append zeros for the non-confounding classes and add them to the original class predictions as $K^{\prime}_{i}=K_{i}+K^{\prime}$. We finally train the linear classifiers $W_{K}$ and $W^{\prime}_{K}$ simultaneously on the CUB train set in a standard supervised setting with cross-entropy loss $L_{ce}$. The total number of classification errors after this class-level intervention procedure is presented in Table 5. We observe a substantial decrease in errors for the confounding classes with an overall increase in the classification performance.
6 Limitations
Despite the numerous advantages, VL-CBMs have two potential limitations.
Ineffable concepts. Though natural language offers a convenient way to build interpretable models with high-level abstract concepts, their expressive power can be limited i.e., the subtle visual cues needed required for solving certain classification tasks, like face recognition, might be difficult to express in words.
Unknown concepts. All CBMs (including VL-CBMs) assume that the complete set of salient concepts required for solving the classification task is known prior to training. However, this requirement cannot be satisfied for all use cases. In such scenarios, prototype-based methods with learnable visual feature prototypes are suitable where they can “discover” unbeknownst concepts during training.
7 Conclusion
In this paper, we investigated the faithfulness of VL-CBM for expert-defined concepts and found that VLMs like CLIP have poor concept alignment. Hence, we proposed a novel Contrastive Semi-Supervised approach that improved the concept alignment in VL-CBM with fewer human-annotated concept labels. We also introduced a class-level intervention procedure for fine-grain classification problems that reduced the error for the confounding classes to increase the overall performance of the model.
8 Acknowledgement
This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of EEE, Nanyang Technological University (NTU), Singapore. This research is supported by DSO National Laboratories, under the project agreement No. DSOCL21238.

### v2

@@
 Abstract
-Concept Bottleneck Models (CBM) map the input image to a high-level human-understandable concept space and then make class predictions based on these concepts.
-Recent approaches automate the construction of CBM by prompting Large Language Models (LLM) to generate text concepts and then use Vision Language Models (VLM) to obtain concept scores to train a CBM.
-However, it is desired to build CBMs with concepts defined by human experts instead of LLM generated concepts to make them more trustworthy.
-In this work, we take a closer inspection on the faithfulness of VLM concept scores for such expert-defined concepts in domains like fine-grain bird species classification and animal classification.
-Our investigations reveal that frozen VLMs, like CLIP, struggle to correctly associate a concept to the corresponding visual input despite achieving a high classification performance.
-To address this, we propose a novel Contrastive Semi-Supervised (CSS) learning method which uses a few labeled concept examples to improve concept alignment (activate truthful visual concepts) in CLIP model.
-Extensive experiments on three benchmark datasets show that our approach substantially increases the concept accuracy and classification accuracy, yet requires only a fraction of the human-annotated concept labels.
-To further improve the classification performance, we also introduce a new class-level intervention procedure for fine-grain classification problems that identifies the confounding classes and intervenes their concept space to reduce errors.
+Concept Bottleneck Models (CBM) map images to human-interpretable concepts before making class predictions.
+Recent approaches automate CBM construction by prompting Large Language Models (LLMs) to generate text concepts and employing Vision Language Models (VLMs) to score these concepts for CBM training.
+However, it is desired to build CBMs with concepts defined by human experts rather than LLM-generated ones to make them more trustworthy.
+In this work, we closely examine the faithfulness of VLM concept scores for such expert-defined concepts in domains like fine-grained bird species and animal classification.
+Our investigations reveal that VLMs like CLIP often struggle to correctly associate a concept with the corresponding visual input, despite achieving a high classification performance.
+This misalignment renders the resulting models difficult to interpret and less reliable.
+To address this issue, we propose a novel Contrastive Semi-Supervised (CSS) learning method that leverages a few labeled concept samples to activate truthful visual concepts and improve concept alignment in the CLIP model.
+Extensive experiments on three benchmark datasets demonstrate that our method significantly enhances both concept ($+29.95\%$) and classification ($+3.84\%$) accuracies yet requires only a fraction of human-annotated concept labels.
+To further improve the classification performance, we introduce a class-level intervention procedure for fine-grained classification problems that identifies the confounding classes and intervenes in their concept space to reduce errors.
 Source codes are available at https://github.com/NMS05/Improving-Concept-Alignment-in-Vision-Language-Concept-Bottleneck-Models.
-Keywords:
-Concept Bottleneck Model Interpretability Semi-Supervised
-1 Introduction
-Concept Bottleneck Models (CBM) [14] are a popular choice for building interpretable models with high-level concepts.
-These models first map the input image to a low-dimensional (hence the term “bottleneck”) expert-defined concept space where each concept is readily understandable by a human.
-For example, the bird species ‘Vermilion Flycatcher’ can have concepts like ‘red-body’, ‘black-wing’, and ‘red-crown’, etc.
-A linear classifier then makes class predictions on top of these concepts.
-There are two fundamental requirements to build a CBM - a well-defined non-ambiguous concept set and the corresponding concept labels for every sample.
-Recent approaches [25, 46] automate the above steps and scale CBM to larger datasets (like ImageNet) by leveraging Large Language Models (LLM) and Vision Language Models (VLM).
-They typically involve generating an initial concept set by prompting a LLM with few shot examples and then extracting a sub-set of those concepts that are conducive for discrimination (classification).
-When an image and this filtered concept set  [^1] The number of LLM generated concepts are usually an order of magnitude larger than the expert-defined concepts.
-is passed as a input to the contrastively pre-trained VLM, it’s image-text alignment scores act as “concept scores or concept labels” for that sample.
-These concept labels can be used to train another CBM classifier [25] or can be directly used to predict the classes [46], thus turning a VLM into CBM (referred to as VL-CBM from here on).
-While LLMs possess rich world knowledge [27, 11, 34] and can generate a potential (large) concept set across any domain, there are certain instances where an expert-defined concept is preferred.
-  • A small and concise set of expert-defined bottleneck concepts enables easier error analysis and intervention, which is crucial for interpretable models.
-  • In medical domains, such as Hematology [36, 2], it is desired to use the expert-defined concept set rather than that generated from LLMs, as the former comes from the more trusted pathologists.
-  • For faithful “visual” classification, all the concepts must be visually groundable, which is guaranteed with expert-defined concept set [38].
-In comparison, for problems like bird species classification, LLMs can generate concepts like “migratory birds” that are not salient for visual classification.
-In such cases with limited yet preferable expert concepts, do the VLM concept scores faithfully represent the visual input and are they efficacious to build VL-CBMs with desired performance?
-To answer this question, we take a closer inspection on the CLIP [28] concept scores for expert-defined concepts on classification datasets like CUB [38], RIVAL [21], and AwA2 [42].
-Ideally, it is desired for the CLIP model to precisely associate a visual concept (appearance, color, shape, etc.) to the corresponding visual input.
-Our investigations expose two problems,
-  • Low Concept Accuracy. The CLIP (VLM) model has a low concept accuracy despite the VL-CBM achieving a high classification performance, which makes the resulting model less faithful.
-  • Incorrect Concept Association.
-For challenging classification problems like CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts.
+Introduction
+Concept Bottleneck Models (CBM) (Koh et al.
+2020) are a popular choice to build interpretable models using human-understandable concepts.
+These models first map input images to a low-dimensional, expert-defined concept space.
+For instance, the bird species ‘Vermilion Flycatcher’ can be described using concepts like ‘red-body’ and ‘black-wing’.
+A linear classifier then makes predictions based on these concepts.
+Building a CBM requires a well-defined, non-ambiguous set of concepts and the corresponding concept labels for each sample, which is labor-intensive to obtain.
+Recent approaches (Oikarinen et al.
+2023; Yang et al.
+2023) circumvent this issue by automating CBM construction using Large Language Models (LLMs) and Vision Language Models (VLMs).
+These methods generate an initial concept set by prompting an LLM with few-shot examples, then extracting a subset of those concepts that are conducive to classification.
+When an image and this filtered concept set are input to a contrastively pre-trained VLM, its image-text alignment scores serve as “concept scores” or “concept labels” for that sample.
+These scores can be used to train another CBM classifier or directly predict the classes, effectively turning a VLM into a CBM (referred to as VL-CBM hereon).
+While LLMs possess rich world knowledge (Petroni et al.
+2019; Jiang et al.
+2020) and can generate extensive concept sets across various domains, expert-defined concepts are sometimes preferred for several reasons.
+  • Concise Concept Sets: A small, well-defined set of expert concepts allows for easier error analysis and intervention, which is crucial for interpretable models.
+  • Trusted Expertise: In medical domains like Hematology (Tsutsui, Pang, and Wen 2023), expert concepts are preferred over LLM-generated ones, as the former comes from the more trusted pathologists.
+  • Visual Grounding: For faithful visual classification, all concepts must be visually grounded in the image, which is ensured with expert concepts.
+In such cases with limited yet preferable expert concepts, it is essential to ask: Do VLM concept scores faithfully represent the visual input, and are they effective for building VL-CBMs with the desired classification performance?
+To answer this, we closely examine CLIP (Radford et al.
+2021) concept scores for expert concepts on classification datasets such as CUB (Wah et al.
+2011), RIVAL (Moayeri et al.
+2022), and AwA2 (Xian et al.
+2018).
+Ideally, the CLIP model should accurately associate visual concepts (appearance, color, shape, etc.) with the corresponding visual input.
+However, our investigations reveal two key issues:
+  • Poor Concept Alignment: The CLIP model exhibits relatively lower concept accuracy despite the higher classification performance of VL-CBMs.
+This suggests that CLIP’s concept scores do not faithfully represent the visual input, thus making the resulting model less reliable.
+  • Challenges with Fine-Grained Concepts: For challenging classification tasks like CUB, the CLIP model tends to be biased towards the primary color of the bird, incorrectly attributing this dominant color to other body parts.
 It struggles to correctly attribute the fine-grain concepts to the visual input.
-Hence, to improve the reliability of VL-CBM, there is a need to improve concept alignment i.e., activate the desired/truthful concepts for a given image, which requires supervision.
-But obtaining the supervisory concept labels for all training samples is cumbersome.
-To solve this problem, we propose a novel Contrastive Semi-Supervised (CSS) learning approach that can improve the concept alignment in VL-CBM with fewer labeled concept examples.
-Our approach encourages consistent concept scores for intra-class samples whilst discriminating (contrasting) them from the other classes.
-It then uses a few labelled concept examples (semi-supervision) per class to align them with the ground truth.
-Extensive experiments on three benchmark datasets CUB [38], RIVAL [21], and AwA2 [42] show that our CSS method substantially increases the concept accuracy (+39.1% for CUB, +18.63% for RIVAL, +31.11% for AwA2) and enhances the overall classification accuracy (+5.61% for CUB, +2.84% for RIVAL, +3.06% for AwA2) with only a handful of human-annotated concept labels per class.
-We also validate our CSS approach in a medical domain using the blood cell attribute dataset WBCAtt [36].
-Intervention is the fundamental advantage of a CBM over other interpretable models - if a classifier makes an incorrect prediction, one can inspect the concept scores and rectify the problem, so that the model makes the desired prediction.
-Intervention in CBM is typically performed at an instance level [14, 25], where a handful of error images are arbitrarily chosen for debugging.
-But, choosing the error images are non-trivial for fine-grain multi-class classification problems which have fewer samples per class and even visually similar classes.
-In this work, we propose a class-level intervention procedure for fine-grain classification problems, where we first compute an error matrix to identify the “confounding classes” (visually similar but semantically different) and then intervene the error images of these classes such that it reduces the total error for the confounding classes whilst increasing the overall classification accuracy.
+Hence, to improve the faithfulness and reliability of VL-CBMs, there is a need to improve concept alignment i.e., activate the desired/truthful concepts for a given image, which requires concept supervision.
+However, obtaining supervisory concept labels for all training samples is cumbersome.
+To address this, we propose a novel Contrastive Semi-Supervised (CSS) learning approach that improves concept alignment in VL-CBM with fewer labeled concept examples.
+Our approach encourages consistent concept scores for intra-class samples while contrasting them with other classes.
+It then aligns these scores with ground truth using a few labeled examples per class (semi-supervision).
+Extensive experiments on three benchmark datasets CUB, RIVAL, and AwA2 show that our CSS method substantially increases the concept accuracy (+39.1% for CUB, +18.63% for RIVAL, +32.11% for AwA2) and enhances the overall classification accuracy (+5.61% for CUB, +2.84% for RIVAL, +3.06% for AwA2) with minimal human-annotated concept labels.
+We also validate the effectiveness our CSS approach in a medical domain using the blood cell attribute dataset WBCAtt (Tsutsui, Pang, and Wen 2023).
+Intervention is the key advantage of a CBM over other interpretable models, where incorrect predictions of the classifier can be rectified by examining and modifying the concepts.
+In this work, we propose a class-level intervention procedure for fine-grain classification problems, where we first compute an error matrix to identify the “confounding classes” (visually similar but semantically different) for which the model makes the most error.
+We then intervene the error images of these classes to improve overall classification accuracy.
 The primary contributions of this paper are summarized below,
-  • Validating the faithfulness of VL-CBM.
-We investigate the concept scores for expert concepts and show that CLIP models have poor concept alignment and they struggle to correctly associate the fine-grain concepts to visual input.
-  • Improving concept alignment.
-To address the above issue and to improve the reliability of VL-CBM, we propose a novel Contrastive Semi-Supervised (CSS) approach which uses a fraction of the human-annotated concept labels to increase the concept accuracy and classification accuracy.
-  • Class-level Intervention.
-We propose an intervention procedure for fine-grain classification problems which reduces the total error for confounding classes and further increases the classification performance.
-2 Related Work
-Post-hoc interpretable models.
-Post-hoc methods [30, 29, 17, 4, 8] convert a pre-trained black-box classifier into an interpretable model using saliency maps.
-They highlight the regions of the input space [31] that influence a classifiers’ prediction and it helps one understand the salient features of a particular class.
-Though saliency maps can reveal the semantic features (or spurious correlations) that a classifier focuses on, their interpretation can be subjective and may require domain expertise.
-Prototype-based and Concept-based interpretable models.
-In stark contrast to post-hoc methods, these models can be trained and are interpretable by design.
-Prototype-based methods [6, 24, 32, 37, 7] learn interpretable visual feature prototypes for every class from the training samples.
-During inference, the distance (similarity score) between the test sample feature and the learned prototypes determine the classifiers decision.
-BotCL [39] is a prototype-based method that is closely related to our work.
-It applies contrastive learning to the prototype space of a vision-only model, while we focus on contrastive + semi-supervised learning on the concept space of a vision-language (multimodal) model.
-Concept-based methods [12, 14, 48] map the input to a high-level concept space defined by human experts, where these concepts are readily interpretable.
-The classifiers decision is based on the predicted concept scores.
-The main advantage of concept-based methods is that they allow test-time intervention [14, 12, 19, 1, 5] and they are easier to debug compared to other methods.
-VLM based interpretable models (VL-CBM).
-The problem with concept-based models is that they require concept labels for every training sample [14] or a lot of training images representing each of those concepts [12, 48].
-To solve this problem, recent approaches [25, 46, 49, 45, 41, 22, 23] leverage contrastively pre-trained Vision-Language Models, like CLIP [28], to automatically generate concept labels from their image-text alignment scores.
-These models have interpretable natural language concepts [18, 44] that are typically generated by a LLM and this automated approach scales them to larger datasets.
-Enhancing faithfulness of VL-CBMs.
-Several recent works have focused on addressing the shortcomings of VLM based concept bottlenecks to improve their reliability.
-They constrain the class predictions to focus on a sub-set of salient concepts [40, 33, 26], make concept scores robust to perturbations [15], and obtain reliable concept score distributions by modeling uncertainties [13] and by composing energy functions [43].
-In our work, we enhance the faithfulness of VL-CBM by using human (expert) preferred concepts and improve the concept scores of VLM in a semi-supervised setup using a few concept labels.
-3 Methodology
-We formally define VL-CBM in Sec.
-3.1 and investigate the faithfulness of CLIP concept scores in Sec.
-3.2.
-We then propose our Contrastive Semi-Supervised (CSS) method to improve concept alignment in Sec.
-3.3.
-3.1 Vision-Language Concept Bottleneck Models
+  • Evaluating faithfulness of VL-CBMs.
+We quantify and evaluate the faithfulness of VLM concept scores, revealing that CLIP models exhibit poor concept alignment and struggle with fine-grain concept association.
+  • Improving Concept Alignment: We introduce a novel Contrastive Semi-Supervised (CSS) approach that leverages a fraction of human-annotated and expert-defined concept labels, which are inherently more reliable, to improve both concept and classification accuracies.
+  • Class-level Intervention: We propose an intervention procedure for fine-grain classification tasks that reduces errors in confounding classes, further enhancing classification performance.
+Related Work
+Early works on Explainable AI focused on unimodal (image) classifiers and they can be broadly classified into two categories.
+Prototype-based methods (Chen et al.
+2019; Nauta, Van Bree, and Seifert 2021; Donnelly, Barnett, and Chen 2022; Wang et al.
+2023) learn interpretable visual feature prototypes for every class from the training samples.
+During inference, the distance (similarity score) between the test sample feature and the learned prototypes determines the classifier’s decision.
+Concept-based methods (Kim et al.
+2018; Koh et al.
+2020; Zarlenga et al.
+2022) map the input to a high-level concept space defined by human experts, where these concepts are readily interpretable.
+The classifier decision is then based on the predicted concept scores.
+The main advantage of concept-based methods is that they allow test-time intervention (Mitchell et al.
+2021; Abid, Yuksekgonul, and Zou 2021; Bontempelli et al.
+2023) and they are easier to debug compared to other methods.
+Vision-Language interpretable models (VL-CBMs).
+A limitation of concept-based models is that they require concept labels for every training sample (Koh et al.
+2020) or a lot of training images representing each of those concepts (Kim et al.
+2018).
+To mitigate this, recent approaches (Oikarinen et al.
+2023; Yang et al.
+2023; Yan et al.
+2023b) leverage contrastively pre-trained Vision-Language Models, like CLIP, to automatically generate concept labels from their image-text alignment scores.
+These models have natural language concepts (Menon and Vondrick 2023; Yan et al.
+2023a) which make them easily interpretable.
+Enhancing faithfulness.
+Though pioneering works on VL-CBMs have proposed methods to extract salient concepts for higher classification accuracy, recent studies have identified several drawbacks, such as concept prompts being biased towards object names (Zang et al.
+2024), concepts being non-factual or irrelevant (Fang et al.
+2024), and insufficient concept coverage (Shang et al.
+2024).
+These recent works improve VL-CBM reliability by discovering concepts that better represent visual attributes (Esfandiarpoor, Menghini, and Bach 2024; Chiquier, Mall, and Vondrick 2024) and have stronger CLIP alignment scores.
+In comparison, we use expert concepts, which are inherently more trustworthy but limited in size.
+Importantly, our CSS method can be applied to these works provided a few labels can be annotated for those concepts.
+Some works also enhance VL-CBMs’ reliability by making the concepts resilient to perturbations (Lai et al.
+2023) and using sparse classifiers (Panousis, Ienco, and Marcos 2023; Semenov et al.
+2024).
+Methodology
+Vision-Language Concept Bottleneck Models
+We first formally define VL-CBMs.
 Consider a dataset $D=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{N},y_{N})\}$, where $x_{i}\in\mathbb{R}^{H\times W\times 3}$ denotes the images, $y_{i}$ denotes the labels for $k$ classes, and $N$ is the total number of training samples.
 While a black-box classifier $f:x\rightarrow\mathbb{R}^{k}$ directly maps an image $x$ to its label $y$, a CBM consists of a backbone $g:x\rightarrow\mathbb{R}^{c}$ which first maps the image $x$ to a concept space consisting of $c$ pre-defined concepts and a linear classifier $h:\mathbb{R}^{c}\rightarrow\mathbb{R}^{k}$ that maps the predicted concepts to its corresponding label $y$.
-These concepts represent the salient visual features that are crucial for solving the classification task.
+These concepts represent the salient visual features crucial for solving the classification task and are human interpretable.
 However, training the backbone $g$ requires $N\times c$ concept labels which is tedious to obtain.
-VL-CBM overcome this issue by leveraging the image-text alignment scores of contrastively pre-trained Vision-Language Models (like CLIP) as concept labels.
+VL-CBMs overcome this issue by leveraging the image-text alignment scores of contrastively pre-trained VLMs (like CLIP) as concept scores.
 Concretely, let $T=\{t_{1},t_{2},...,t_{c}\}$ be the set of LLM generated or expert-defined text concepts.
-Now the $c$ concept labels for an image $x_{i}$, which is denoted as $C_{i}$, is obtained by $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}$, where $\cdot$ denotes the dot product, $E_{I}$ and $E_{T}$ denotes the image and text encoder of the CLIP model that maps the images and text concepts into a shared feature space.
-Now, with the availability of image $x_{i}$, concept label $C_{i}$, and class label $y_{i}$ for $N$ samples, the backbone $g$ and the classifier $h$ can be independently or jointly trained.
-Table 1: Evaluation of CLIP concept scores and VL-CBM classification performance.
-   CUB    RIVAL    AwA2 Classification Accuracy (%)    75.84    95.63    90.14 Concept Accuracy (%)    24.43    58.85    49.02 No.of Concepts    312    18    85 Difficulty of Concepts    Hard    Easy    Medium
-3.2 Faithfulness of Vision-Language Concept Bottleneck Models
-We validate the faithfulness of VL-CBM for expert concepts across three image classification datasets namely Caltech-UCSD Birds (CUB) [38], Rich Visual Attributes with Localization (RIVAL) [21], and Animal with Attributes (AwA2) [42].
-These datasets provide expert concepts and the corresponding ground-truth concept labels required for our evaluation.
-We input the images and the expert (text) concepts to a frozen CLIP model to obtain the concept scores as described in Sec.
-3.1.
-Our objective here is to evaluate the quality and faithfulness of the VLM concept scores by comparing them against the ground-truth concept labels and evaluate their classification performance with a linear classifier.
-The concept evaluation results are presented in Table 1.
-Our investigations expose two problems with CLIP concept scores.
-Low Concept Accuracy.
-From the table, the CLIP model has a low concept accuracy despite the VL-CBM achieving a high classification accuracy.
-This implies that a higher classification performance does not guarantee a faithful concept understanding, which makes the VL-CBM less reliable.
-Moreover, it becomes difficult to intervene such a model during test-time, in order to rectify prediction errors.
-Incorrect Concept Association.
-A closer inspection of these concepts in Fig.
-2 (CLIP Concept Scores) reveal that, for fine-grained concepts like in CUB, the CLIP model is sensitive and biased towards the primary color of the bird and associate this dominant color to all other body parts.
-The untrained model struggles to correctly associate the color/texture concepts to the corresponding bird part, which is crucial for bird species classification.
-The lower concept accuracy and the errors in concept association is a result of the poor relational and compositional understanding of CLIP models i.e., they make errors in associating objects to their attributes [47], especially in a multi-object and multi-attribute setting [16].
-Hence, to improve the reliability and faithfulness of VL-CBM, there is a need to improve concept alignment i.e.
-activate the truthful concepts for the given image.
-Figure 1: (a) Architecture of VL-CBM with concept projection (bootstrap). (b) Overview of the Contrastive Semi-Supervised (CSS) learning method.
-3.3 Improving concept alignment in VL-CBM
-We propose a novel Contrastive Semi-Supervised (CSS) learning method that can improve the concept alignment in VL-CBM with a few labelled concept examples per class.
-The overall framework is illustrated in Fig.
-1.
-From Table 1, we can see that the frozen CLIP model achieves a low yet non-trivial concept accuracy.
-We bootstrap this by adding a learnable linear concept projection layer (see Fig.
-1a) that enhances the CLIP concept scores by directly predicting the $c$ concepts from the average pooled visual patch tokens as,
+Now the $c$ concept scores for an image $x_{i}$, which is denoted as $C_{i}$, is obtained by $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}$, where $\cdot$ denotes the dot product, $E_{I}$ and $E_{T}$ denotes the image and text encoder of the CLIP model that maps the images and text concepts into a shared feature space.
+Now, with the availability of image $x_{i}$, concept score $C_{i}$, and class label $y_{i}$ for $N$ samples, the linear classifier $h$ can be trained to make class predictions.
+   CUB    RIVAL    AwA2 Classification Acc (%)   75.84   95.63   90.14 Concept Acc (%)    24.43    58.85    49.02 No.of Concepts    312    18    85 Concept Difficulty    Hard    Easy    Medium
+Table 1: Evaluating faithfulness of CLIP concept scores and VL-CBM classification performance.
+Evaluating Faithfulness of VL-CBMs
+The trustworthiness and reliability of a VL-CBM depends on the faithfulness of its concept scores, which serve two purposes: providing rich representations for a linear classifier and enabling human interpretation of the classifier’s predictions.
+Prior works on VL-CBMs have focused primarily on classification accuracy, which measures how well samples are separable in the concept space, but doesn’t measure the faithfulness of concept scores in representing visual truth.
+Therefore, we evaluate both the faithfulness of VLM concept scores by comparing them to ground-truth concept labels (Koh et al.
+2020) and their classification performance with a linear classifier.
+We use three image classification datasets - Caltech-UCSD Birds (CUB) (Wah et al.
+2011), Rich Visual Attributes with Localization (RIVAL) (Moayeri et al.
+2022), and Animal with Attributes (AwA2) (Xian et al.
+2018) because they provide expert-defined concepts and the corresponding ground-truth concept labels needed for our evaluation.
+The results are presented in Table 1 and a visualization of these concepts is shown in Figure 2.
+Our investigations reveal two key issues with CLIP concept scores.
+Poor Concept Alignment: From the table, the CLIP model has relatively lower concept accuracy across all three datasets, despite the VL-CBM achieving high classification accuracy.
+This indicates that CLIP’s concept scores do not faithfully represent the visual input and a higher classification performance does not guarantee a faithful concept understanding.
+Consequently, the model is less reliable, and it becomes difficult to intervene such a model during test-time.
+Challenges with Fine-Grained Concepts: A closer inspection of these concepts in Figure 2 (CLIP Concept Scores) reveals that, for challenging tasks like CUB, the CLIP model tends to be biased towards the primary color of the bird and incorrectly attributes this dominant color to other body parts.
+It struggles to correctly associate the compositional and fine-grain concepts with the visual input.
+Why does CLIP exhibit poor concept alignment?
+Although the CLIP model excels at associating images with text containing object names and attributes, achieving strong zero-shot performance on open-ended classification tasks, we find that the CLIP model struggles with fine-grained visual understanding when only concept descriptions are presented, making errors in associating objects with their attributes.
+The CLIP model shows unsatisfactory concept accuracy when dealing with such fine-grained concepts defined by experts.
+Similar findings are also discussed in some previous works (Zang et al.
+2024; Pham et al.
+2024; Yuksekgonul et al.
+2022; Lewis et al.
+2022; Yun et al.
+2022).
+As poor concept alignment can reduce the reliability and interpretability of VL-CBMs, it is critical to improve the faithfulness of VL-CBMs by activating the truthful concepts for the given image.
+Figure 1: (a) Modified CLIP (VLM) with concept projection layer which enhances the raw concept scores.
+(b) Architecture of VL-CBM and an overview of the Contrastive Semi-Supervised (CSS) learning method.
+Enhancing Faithfulness of VL-CBMs
+To improve concept alignment in VL-CBM we propose a novel Contrastive Semi-Supervised (CSS) learning method that leverages a few labelled concept examples per class.
+Figure 1 illustrates the overall framework.
+As shown in Table 1, the frozen CLIP model achieves a low yet non-trivial concept accuracy.
+We enhance this by adding a learnable linear concept projection layer (see Figure 1a) which improves the CLIP concept scores by directly predicting $c$ concepts from the average pooled visual patch tokens as,
 $C_{i}=E_{I}(x_{i})\cdot E_{T}(T)^{\intercal}+LN(E_{I}^{*}(x_{i}))\cdot W_{CP},$ (1)
@@
 From the concept scores $C_{i}\in\mathbb{R}^{1\times c}$, a linear classifier $W_{K}\in\mathbb{R}^{c\times k}$ makes class predictions $K_{i}=C_{i}\cdot W_{K}$ for the image $x_{i}$, where $K_{i}\in\mathbb{R}^{1\times k}$ denotes the class logits for the $k$ classes.
-Now the model must be optimized to improve concept alignment and achieve high classification accuracy.
-To improve concept alignment, the predicted concepts can be compared with the ground truth concept labels.
-However, the availability of fewer concept labels (semi-supervision) makes it less effective in improving concept alignment across other unlabelled samples.
-To mitigate this, we augment a contrastive learning objective to the concept space with the following intuition.
-The contrastive objective encourages “consistent” concept scores for intra-class samples and the concept labels for a few of these samples (semi-supervision) guide or align the predicted concepts with the ground truth.
-Furthermore, contrastive learning discriminates the concept scores of inter-class samples thus aiding in classification.
-Such a learning objective of enforcing consistency between sample pairs guided by a few labeled examples [3, 20] has not been studied in VL-CBMs before.
-To achieve this (see Fig.
-1b), we perform pairwise sampling such that each training sample consists of a pair of images $(x_{i},x_{j})$ from the same class and every other sample in the mini-batch belongs to different classes.
-There are $n$ samples ($2n$ images) in each mini-batch.
+Now, the model must be optimized to enhance both concept and classification accuracies.
+To improve concept accuracy, the concept projection layer can be trained with a few concept labels (semi-supervision), which refines the weak concept scores of CLIP and aligns the predicted concepts with the ground truth.
+However, one needs to effectively leverage this sparse (concept) supervisory signals and extend them to other unlabelled examples.
+Contemporary unimodal works (Hu et al.
+2024) compare the features of labelled and unlabelled examples in the cosine space, and then assign pseudo concept labels for unlabelled examples using KNN algorithm.
+But, this necessitates additional computation steps (KNN) for each mini-batch.
+We instead propose to bootstrap the sparsely supervised concept scores to other unlabelled samples using contrastive learning.
+This method promotes consistent concept scores for samples sharing ‘similar’ concepts (positive pairs) and discriminates concept scores for samples with ‘disjoint’ concepts (negative pairs).
+For all datasets in our paper, the similar concept pairs come from intra-class samples and the disjoint concept pairs come from inter-class samples.
+Furthermore, discriminating the scores of inter-class samples make classes more separable in the concept space, thus aiding in classification.
+To achieve this (see Figure 1b), we perform pairwise sampling such that each training sample consists of a pair of images $(x_{i},x_{j})$ from the same class, while every other sample in the mini-batch belongs to different classes.
+There are $n$ samples ($2n$ images) in each mini-batch, and only a subset of these images will be concept supervised.
 For these image pairs, we compute the concept scores $(C_{i},C_{j})$ and class predictions $K=[K_{i},K_{j}]~{}$ [^1] The square brackets denote the tensors are stacked across batch dimension and no longer need to be in pairs..
@@
 We jointly train the concept projection layer and the linear classifier to minimize the following loss objective.
-[…]
 $L=L_{contrastive}+L_{CE}+L_{concept},$ (2)
@@
 $L_{concept}=0,\ \ \textrm{if}\ \ |G_{l}|=0,$
-where $sim$ is the cosine similarity metric, $\tau$ is the temperature parameter, $L_{ce}$ and $L1$ are the standard cross entropy loss and mean absolute error, $s$ is the softmax normalization and $\gamma$ is a constant that increases the magnitude of the $L1$ loss function.
+where $sim$ is the cosine similarity metric, $\tau$ is the temperature parameter, $L_{ce}$ and $L1$ is the standard cross entropy loss and mean absolute error, $s$ is the softmax normalization and $\gamma$ is a constant that increases the magnitude of the $L1$ loss function.
+$G_{l}$ represents the ground truth concept label for the $l$-th entry.
 $l$ denotes the index of the concept score, class logit, and label tensor within the mini-batch, after stacking along the batch dimension.
@@
 The concept loss is enabled if the concept labels are available ($|G_{l}|\neq 0$) or else it becomes zero.
-4 Experiments
-4.1 Experimental Setup
-Dataset.
-We validate the effectiveness of our Contrastive Semi-Supervised (CSS) method in improving concept alignment on the CUB [38], RIVAL [21], AwA2 [42] datasets, and in a medical setting on the White Blood Cell Attribute (WBCAtt) dataset [36].
-All the above datasets provide an expert-defined concept set and the corresponding ground-truth concept labels, which is crucial for our evaluation.
-Model and Training.
-We use the PLIP ViT-B-16 [9] model available in Hugging Face for experiments on WBCAtt dataset and use the OpenCLIP ViT-B-16 model [10] for experiments on all other datasets.
-We use Adam optimizer to train all our models and evaluate them with classification accuracy and/or concept accuracy metrics.
+Model CUB RIVAL AwA2 WBCAtt  Class  Concept  Class  Concept  Class  Concept Attribute ViT-B/16 (Black Box)  79.21 NA  99.47 NA   94 NA  66.11 CLIP VL-CBM  75.84 24.43  95.63 58.85  90.14 49.02 42.56 CSS VL-CBM (Ours)  81.45 63.53  98.47 77.48   93.2  81.13  64.19
 Table 2: Improving Concept Alignment in VL-CBM using Contrastive Semi-Supervised (CSS) learning.
@@
 Note that WBCAtt uses the PLIP model.
-Model CUB RIVAL AwA2 WBCAtt  Class  Concept  Class  Concept  Class  Concept Attribute ViT-B/16 (Black Box)  79.21 NA  99.47 NA   94 NA  66.11 CLIP VL-CBM [25, 46]  75.84 24.43  95.63 58.85  90.14 49.02 42.56 CSS VL-CBM (Ours)  81.45 63.53  98.47 77.48   93.2  81.13  64.19
-4.2 Improving Concept Alignment
+Figure 2: Visualization of $top-8$ concept scores for the CUB (birds) and AwA2 (animals) datasets. Incorrectly activated concepts are highlighted in red.
+Experiments
+Experimental Setting
+Dataset.
+We validate the effectiveness of our Contrastive Semi-Supervised (CSS) method in improving concept alignment on the CUB (Wah et al.
+2011), RIVAL (Moayeri et al.
+2022), AwA2 (Xian et al.
+2018) datasets, and in a medical setting on the White Blood Cell Attribute (WBCAtt) dataset (Tsutsui, Pang, and Wen 2023).
+All the above datasets provide an expert-defined concept set and the corresponding ground-truth concept labels, which is crucial for our evaluation.
+Model.
+We use the PLIP ViT-B-16 (Huang et al.
+2023) model available in Hugging Face for experiments on WBCAtt dataset and use the OpenCLIP ViT-B-16 model (Ilharco et al.
+2021) for experiments on all other datasets, unless otherwise specified.
+Training and Evaluation.
+Following the literature (Koh et al.
+2020; Fong and Vedaldi 2017), we use concept accuracy to quantify concept alignment.
+We use Adam optimizer to train all our models and evaluate them with classification accuracy and concept accuracy metrics.
+We also conduct a distributional analysis of the VL-CBM concept scores to show the concept alignment explicitly.
+Improving Concept Alignment
 We empirically demonstrate the effectiveness of our Contrastive Semi-Supervised (CSS) approach in improving concept alignment for the CUB, RIVAL, AwA2, and WBCAtt datasets in Table 2.
 Bird and Animal classification.
-We validate our CSS method with concept accuracy and classification accuracy metrics for CUB, RIVAL, and AwA2 datasets.
-We first train a black-box image classifier with frozen CLIP vision encoder (ViT-B/16) as the backbone to establish the baseline classification performance.
-The black-box classifier directly predicts the classes from the image features without any concept bottlenecks.
-We then train a CLIP VL-CBM classifier with expert-defined text concepts as described in Sec.
-3.1.
-From the table, we can see that the classification performance of CLIP VL-CBM is on-par with the black-box classifier, however, it has a lower concept accuracy.
-Figure 2: Visualization of $top-8$ concept scores for the CUB (birds) and AwA2 (animals) datasets. Incorrectly activated concepts are highlighted in red.
-Hence, to improve the faithfulness and reliability of VL-CBM, we apply the concept projection later (bootstrap) and optimize the model with our CSS objective (CSS VL-CBM).
+We validate our CSS method with concept accuracy and classification accuracy metrics for the CUB, RIVAL, and AwA2 datasets.
+We first train a black-box image classifier with frozen CLIP vision encoder (ViT-B/16) as the backbone to establish baseline classification performance.
+The black-box classifier directly predicts the classes from image features without any concept bottlenecks.
+We then train a CLIP VL-CBM (Oikarinen et al.
+2023; Wang et al.
+2023) classifier with expert-defined text concepts.
+From the table, we can see that the classification performance of CLIP VL-CBM is on par with the black-box classifier, however, it has a lower concept accuracy.
+Hence, to improve the faithfulness of VL-CBM, we add the concept projection layer and optimize the model with our CSS objective (CSS VL-CBM).
 Our CSS method not only substantially increases the concept accuracy, it also closes the gap with black-box methods on image classification performance.
@@
 The results show that even with such small percentages of concept labels, our CSS method can substantially increase the concept accuracy.
-The enhanced and faithful concept scores after improving concept alignment with CSS is presented in Fig.
-2 (CSS Concept Scores).
+The enhanced and faithful concept scores after improving concept alignment with CSS is presented in Figure 2 (CSS Concept Scores).
 WBC Attribute prediction.
-WBCAtt contains White Blood Cell (WBC) images annotated with 31 “cell morphology” attributes that aids in diagnosis in Hematology.
-We directly evaluate the models on their attribute prediction accuracy and do not perform any image classification.
-While contemporary approaches [48, 46] use the HAM10000 [35] dataset to study VL-CBM in the medical domain, the availability of attribute labels in WBCAtt makes the latter suitable for our work.
-For our evaluation, we train the PLIP model [9] (a VLM model for Pathology AI), with 50% of the attribute labels.
+WBCAtt contains White Blood Cell (WBC) images annotated with 31 “cell morphology” attributes.
+These attributes aid in diagnosis of a variety of conditions in Hematology and Pathology.
+So, we directly evaluate the models on their attribute prediction accuracy and do not perform any image classification.
+While contemporary works (Yuksekgonul, Wang, and Zou 2023; Yang et al.
+2023) use the HAM10000 (Tschandl, Rosendahl, and Kittler 2018) dataset to study VL-CBM in the medical domain, the availability of attribute labels in WBCAtt makes the latter suitable for our work.
+For our evaluation, we train the PLIP model (a VLM for Pathology AI), with 50% of the attribute labels.
 From the table, we can see that our CSS method substantially increases the attribute prediction accuracy and closes the gap with black-box attribute predictor.
-Figure 3: Number of concept labels vs CSS VL-CBM performance.
+Figure 3: Impact of the number of concept labels on CSS VL-CBM and LaBo.
 Impact of the number of concept labels.
-Since our approach is semi-supervised, we further investigate the impact of the number of human-annotated concept labels on the CSS VL-CBM performance in Fig.
-3.
-For CUB and RIVAL datasets we progressively increase the concept labels used for training, for which we observe a proportional increase in the classification and concept accuracies.
-For WBCAtt, we investigate how the black-box (PLIP vision encoder) attribute predictor compares against the CSS (PLIP) VL-CBM attribute predictor, for an increase in the attribute labels used for training.
-Note that there is no image classification task for WBCAtt.
-We can see that our CSS method outperforms black-box attribute predictor in a low data-resource setting.
-4.3 Comparison with other methods
+Since our approach is semi-supervised, we further investigate the impact of the number of human-annotated concept labels on CSS VL-CBM performance, as shown in Figure 3.
+In this study, we also compare our method with the closest work, LaBo, where we fine-tune the final layernorm and visual projection layers of the CLIP model (which closely resembles training our concept projection layers).
+Both models are trained with expert concepts, and we progressively increase the number of concept labels used for training on the CUB and RIVAL datasets.
+From the figure, we observe that both approaches achieve similar classification accuracies.
+However, our CSS VL-CBM model attains relatively higher concept accuracies, especially when fewer concept labels are available.
+We attribute this to the use of the concept projection layer (which enhances raw CLIP scores) and the application of contrastive learning.
+CBM Method Concepts Source Class Acc (%) LaBo (Yang et al.
+2023) 10,000 LLM 81.9 SparseCBM (Semenov et al.
+2024) 926 LLM 80.02 CDL (Zang et al.
+2024) 400 LLM 83.4 LM4CV (Yan et al.
+2023a) 400 LLM 81.4 CSS VL-CBM (Ours) 321 Expert 83.89
+Table 3: Comparison of CSS VL-CBM with contemporary methods on CUB dataset.
+Comparison with other methods
 CUB.
-The CUB dataset is extensively used by the Explainable AI community to build and study concept-based interpretable models.
-Hence, we comprehensively compare our Contrastive Semi-Supervised VL-CBM method with contemporary VLM-based interpretable models on the CUB dataset.
-The results are presented in Table 3.
-For a fair comparison with LF-CBM [25] and PCBM [48], we perform two-stage training.
-In stage 1, a ResNet18 backbone with a linear concept predictor is trained with the concept labels generated by our CSS VL-CBM model.
-It minimizes the $L_{concept}$ loss function.
-In stage 2, a sparse-linear SGD classifier from scikit-learn is trained with 5000 steps as described in [48].
-The backbone, concept predictor, and sparse-linear classifier together constitute a CBM.
-Table 3: Comparison of CSS VL-CBM with contemporary methods on CUB.
-Method Base CBM Training Classifier Class (%) Concept (%) LF-CBM [25] ResNet-18 2-stage Sparse-Linear 74.31 - PCBM [48] ResNet-18 2-stage Sparse-Linear 59.6 - CLIP VL-CBM ResNet-18 2-stage Sparse-Linear 68.77 16.53 CSS VL-CBM (Ours) ResNet-18 2-stage Sparse-Linear 72.85 62.8 CLIP VL-CBM ResNet-18 joint Linear 74.91 18.84 CSS VL-CBM (Ours) ResNet-18 joint Linear 80.36 63.17 CompDL [49] CLIP ViT-B/32 - Linear 52.6 - LaBo (full shot) [46] CLIP ViT-B/32 - Linear 81.0 - CSS VL-CBM (Ours) CLIP ViT-B/16 joint Linear 81.45 63.53
-From the table, we can see that the classification accuracy of ResNet-18 CBM supervised by our CSS VL-CBM approach is on par with LF-CBM and achieves much higher performance than PCBM.
-However, we observe the best performance with joint training of the backbone, concept predictor, and linear classifier layers with $L_{concept}$ and $L_{ce}$ losses.
-Grad-CAM [31] visualization of the concepts learned by the (jointly trained) ResNet18 CBM supervised by our CSS VL-CBM is presented in Fig.
-4.
-From the figure, we can see that the improved concept knowledge (activating the truthful concepts for a given image) of our CSS VL-CBM model was effectively transferred to the ResNet model.
-It is able to correctly associate a color or texture concept to the corresponding bird part.
-Also, our method performs better than LaBo [46] with just $312$ (CUB) expert concepts, while LaBo utilizes $10,000$ LLM-generated bottleneck concepts.
-Figure 4: Grad-CAM visualization of the learned concepts.
-Table 4: Comparison of CSS VL-CBM with other methods on RIVAL and AwA2.
-Dataset RIVAL AwA2 CBM Method  Text2Concept   CSS  PCBM   ECBM   CSS Classification Accuracy (%) 95.38   98.86  88   91.2   93.2 Concept Accuracy (%) -   72.84  71.9   85.84   81.3
+We compare our Contrastive Semi-Supervised VL-CBM method with contemporary VLM-based interpretable models on the CUB dataset, which is extensively used by the Explainable AI community to study CBMs.
+Following the literature, we adopt the CLIP ViT-L-14 model for this comparison.
+Note that this increases the number of trainable parameters in the concept projection layer.
+The results, presented in Table 3, show that our CSS model achieves strong classification performance despite having a smaller number of bottleneck concepts.
+Since other methods use LLM-generated concepts without corresponding concept labels, concept accuracy is not reported for this comparison.
+CBM Method Class Acc (%) Concept Acc (%) Text2Concept 95.38 _ CSS (Ours) 98.86 72.84 PCBM 88 71.9 ECBM 91.2 85.84 CSS (Ours) 93.2 81.3
+Table 4: Comparison of CSS VL-CBM with other methods on RIVAL (top) and AwA2 (bottom) datasets.
 RIVAL and AwA2.
 We further compare CSS VL-CBM with contemporary methods on the animal classification datasets in Table 4.
-For RIVAL dataset, CSS achieves a higher classification performance compared to Text2Concept [22].
-For AwA2 dataset, CSS achieves a higher classification accuracy compared to other methods, but achieves a slightly lower concept accuracy compared to ECBM.
-It is important to note that PCBM [13] and ECBM [43] do not use “text” concepts and they optimize the feature extractor (vision backbone) during training.
-5 Intervention
-When a CBM makes an error, one can inspect the incorrect predictions and manually tune the classifier weights or modify the concepts to fix those errors.
-This approach is called “intervention” or “model editing”.
-Conventional approaches to intervention are typically done at the instance-level [25, 48, 14], where a handful of error images are arbitrarily chosen for intervention.
-However, for fine-grain multi-class classification problems like CUB which have more classes (some bird species are visually similar) and relatively fewer samples per class, choosing the error images is non-trivial.
-Moreover, it is desired to perform intervention such that it increases the performance globally [48] instead of rectifying the errors for a few images.
-Hence, we propose a class-level intervention procedure where we first identify the confounding classes and then choose the error images of these classes for intervention.
-Confounding classes are bird species, that are visually similar yet semantically different and the model makes the most errors for these challenging classes.
-This class-level intervention is done in the following steps.
-Figure 5: Confounding classes - birds that are visually similar yet belong to different sub-species.
+For the RIVAL dataset, CSS achieves higher classification performance compared to Text2Concept (Moayeri et al.
+2023).
+For the AwA2 dataset, CSS achieves higher classification accuracy compared to other methods but slightly lower concept accuracy compared to ECBM.
+It is important to note that PCBM (Kim et al.
+2023) and ECBM (Xu et al.
+2023) do not use “text” concepts and optimize the vision backbone during training.
+Metric CSS Scores CLIP Scores Truthfulness ($\downarrow$) 3.22 14.04 Sparseness# ($\downarrow$) 0.08 0.19 Discriminability# ($\uparrow$) 7.04 3.28
+Table 5: Distributional analysis of CSS vs CLIP concept scores for CUB test. The results are averaged over instances and classes#.
+Distributional Analysis of Concept Space
+In addition to concept accuracy metric, we conduct a distributional analysis of the VL-CBM concept scores to explicitly demonstrate the effectiveness of our CSS method in improving concept alignment.
+For quantitative assessment, we extract the concept scores ($\mathbb{R}^{5790\times 312}$) from both CLIP and our CSS model for the CUB test set.
+We report three metrics: truthfulness: defined as the L2-distance between predicted and ground-truth concepts; sparseness: measures the intra-class standard deviation of the predicted concepts; and discriminability: defined as the L2-distance between inter-class concept score clusters.
+The results in Table 5 show that the proposed CSS model has improved concept alignment compared with the CLIP model.
+To further visually compare the difference, Figure 4 depicts a t-SNE visualization of CSS (triangles) vs.
+CLIP (circles) concept scores of 5 random classes on the CUB test set.
+Our CSS method makes the concept clusters more dense and more separable in the concept space, enhancing concept alignment by bringing them closer to the ground truth.
+Figure 4: t-SNE visualization of CSS (triangles) vs CLIP (circles) concept scores of 5 random classes for CUB test.
+Intervention
+When a CBM makes an error, one can inspect the incorrect predictions and manually adjust the classifier weights or modify the concepts to rectify these errors, thereby improving the overall accuracy.
+This process is known as “intervention” or “model editing.” Traditional intervention approaches are typically performed at the instance level (Oikarinen et al.
+2023; Yuksekgonul, Wang, and Zou 2023; Koh et al.
+2020), where a few erroneous images are arbitrarily chosen for intervention.
+However, for fine-grained multi-class classification problems like CUB, which have many classes (some bird species are visually similar) and relatively few samples per class, selection of error images become non-trivial.
+Therefore, we propose a class-level intervention procedure where we first identify the confounding classes and then choose the error images from these classes for intervention.
+Confounding classes are bird species that are visually similar yet semantically different, and the model makes the most errors for these challenging classes.
+This class-level intervention is carried out in the following steps.
+Step 1: Identify the confounding classes.
+We compute a $K\times K$ error matrix by evaluating the trained CSS VL-CBM model on the test set, where $K$ is the number of classes.
+The error matrix provides the error distribution for each class, and the matrix element with the highest value indicates the confounding class pairs.
+In our experiments, we identified two pairs of confounding classes: California Gull - Western Gull and Common Tern - Arctic Tern.
+The total number of classification errors for these confounding classes before intervention is presented in Table 6.
+Step 2: Inspecting the concept space.
+Next, we inspect the concept scores of the error images in these confounding class pairs.
+We find that the concept score distribution of the confounding classes is nearly identical, making it difficult to discriminate between them.
+Even the averaged ground-truth concept labels for the confounding classes are highly similar.
+Our model suffers from Type-2 error (Oikarinen et al.
+2023), meaning the number of bottleneck concepts and their score distributions are insufficient to effectively discriminate the samples.
+Before Interv.
+After Interv.
+Total error for CG (30) 27 14 Total error for WG (30) 15 9 CG misclassified as WG 13 4 WG misclassified as CG 3 1 Total error for CT (30) 21 11 Total error for AT (29) 3 2 CT misclassified as AT 12 6 AT misclassified as CT 0 0 Classification Accuracy (%) 81.45 82.57
 Table 6: Class-level intervention of CSS VL-CBM.
@@
 The value within brackets denote the total number of test samples for that class.
-Before Interv.
-After Interv.
-Total error for CG (30) 27 14 Total error for WG (30) 15 9 CG misclassified as WG 13 4 WG misclassified as CG 3 1 Total error for CT (30) 21 11 Total error for AT (29) 3 2 CT misclassified as AT 12 6 AT misclassified as CT 0 0 Overall Classification Accuracy (%) 81.45 82.57
-Step 1: Identify the confounding classes.
-First, we compute a $K\times K$ error matrix by evaluating the trained CSS VL-CBM model on the test set, where $K$ is the number of classes.
-The error matrix gives the error distribution for every class and the index of the matrix element with the highest value (error) gives the confounding class pairs.
-In our experiments, we identified two pairs of confounding classes “California Gull - Western Gull” and “Common Tern - Artic Tern” as shown in Fig.
-5.
-The total number of classification errors for these confounding classes before intervention are presented in Table 5.
-Step 2: Inspecting the concept space.
-Next, we inspect the concept scores of the error images in these confounding class pairs.
-We find that the concept score distribution of the confounding classes are nearly identical, which makes it difficult to discriminate them.
-Our model suffers from Type-2 error [25] i.e., the number of bottleneck concepts and their score distributions are insufficient to effectively discriminate the samples.
 Step 3: Expanding the concept set.
-To mitigate the Type-2 error and to improve the overall performance, we expand the concept set by leveraging the concepts provided by LaBo [46].
+To mitigate the Type-2 error and to improve the overall performance, we expand the concept set by leveraging the concepts provided by LaBo (Yang et al.
+2023).
 Note that the new concepts to be added should be visually groundable by the CLIP model and hence we use the “filtered” concept set of LaBo.
@@
 We observe a substantial decrease in errors for the confounding classes with an overall increase in the classification performance.
-6 Limitations
-Despite the numerous advantages, VL-CBMs have two potential limitations.
+Limitations
+Despite the numerous advantages, VL-CBMs have a few potential limitations.
 Ineffable concepts.
-Though natural language offers a convenient way to build interpretable models with high-level abstract concepts, their expressive power can be limited i.e., the subtle visual cues needed required for solving certain classification tasks, like face recognition, might be difficult to express in words.
+While natural language offers a convenient way to build interpretable models with high-level abstract concepts, its expressive power can be limited.
+The subtle visual cues required for certain classification tasks, such as face recognition, might be difficult to articulate in words.
 Unknown concepts.
-All CBMs (including VL-CBMs) assume that the complete set of salient concepts required for solving the classification task is known prior to training.
-However, this requirement cannot be satisfied for all use cases.
-In such scenarios, prototype-based methods with learnable visual feature prototypes are suitable where they can “discover” unbeknownst concepts during training.
-7 Conclusion
-In this paper, we investigated the faithfulness of VL-CBM for expert-defined concepts and found that VLMs like CLIP have poor concept alignment.
-Hence, we proposed a novel Contrastive Semi-Supervised approach that improved the concept alignment in VL-CBM with fewer human-annotated concept labels.
+VL-CBMs assume that the complete set of salient concepts required for solving the classification task is known before training.
+However, this assumption is not always feasible.
+In such cases, models that can “discover” unknown concepts (Shang et al.
+2024) during the course of training are more suitable.
+Locality Faithfulness.
+Recent works have observed that even when CBMs predict the concept correctly, they often fail to ‘respect localities’ (Raman et al.
+2024).
+This suggests that the concept predictor might rely on spurious correlations instead of relevant visual/spatial features.
+Such a study on locality faithfulness in VL-CBMs is largely unexplored.
+Conclusion
+In this paper, we investigated the faithfulness of VL-CBM for expert concepts and found that VLMs like CLIP have poor concept alignment.
+Hence, we proposed a novel Contrastive Semi-Supervised approach that improved concept alignment with fewer human-annotated concept labels.
+Our work will primarily aid the Explainable AI community in building interpretable models that align with human preferences (concepts) yet require minimal supervision.
 We also introduced a class-level intervention procedure for fine-grain classification problems that reduced the error for the confounding classes to increase the overall performance of the model.
-8 Acknowledgement
-This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of EEE, Nanyang Technological University (NTU), Singapore.
-This research is supported by DSO National Laboratories, under the project agreement No.
-DSOCL21238.